# Entanglement, Spin Squeezing and Quantum Sensing in a Spin-5/2 Heisenberg Molecular Iron(III) Triangle

### v2

@@ Abstract
This study provides insights into the static and dynamic quantum properties of the trinuclear high-spin iron(III) molecular complex $[\mathrm{Fe}_{3}\mathrm{Cl}_{3}(\mathrm{saltag^{\mathrm{Br}}})(\mathrm{py})_{6}]\mathrm{ClO}_{4}$ to be further abbreviated as Fe3. Using exact diagonalization of a spin-5/2 Heisenberg triangle in a magnetic field, we model the corresponding quantum behavior of the molecular compound Fe3. Our rigorous analysis employs various key metrics to explore a rich quantum behavior of this molecular compound. At sufficiently low temperatures, the bipartite negativity reveals that the pairwise entanglement between any pair of iron(III) magnetic ions of the molecular complex Fe3 can be significantly enhanced by a small magnetic field. This enhancement is followed by unconventional step-like changes characterized by a sequence of plateaus and sudden downturns as the magnetic field further increases. A qualitatively similar behavior is also observed in the genuine tripartite entanglement among all three iron(III) magnetic ions in the trinuclear complex Fe3. Notably, the bipartite and tripartite entanglement persist in the molecular complex Fe3 up to moderate temperatures of approximately 30 K and 70 K, respectively. We also utilized a spin squeezing parameter to identify parameter regions of coherent, entangled, and squeezed states of the molecular compound Fe3. Our findings demonstrate that the squeezed states, which are relevant for technological applications, can be achieved at temperatures below 15 K and magnetic fields under 25 T. Additionally, we demonstrate the achievement of quantum-enhanced sensitivity by initializing the molecular complex Fe3 in Dicke states. Finally, we investigated a quantum-sensing protocol by applying a local magnetic field specifically to one iron(III) magnetic ion of the molecular compound Fe3 and performing readout sequentially on one of two remaining iron(III) magnetic ions.
@@ I Introduction
The field of molecular magnetism Kahn ; Georges2001 ; Schnack2004 ; Schnack2005 is rapidly advancing with combining insights from quantum information science, condensed matter physics and chemical engineering Coronado2020 ; Pineda2021 ; Leuenberger2001 ; Chiesa2024 . Molecular magnets, with their unique features, have promising applications in ultra-high-density storage devices Qin2009 ; Baltic2016 , spintronics Sanvito2011 ; Fursina2023 , quantum computing Fursina2023 ; Song2014 ; Kandala2017 ; Carretta2021 ; Ollitrault2021 ; Giansiracusa2021 ; Chizzini2022 , and quantum algorithms Godfrin2017 . Phenomena like quantum tunneling of magnetization and molecular cooling in the quantum magnets are essential for quantum computing and achieving low temperatures Thomas1996 ; Gatteschi2003b ; Wu2021 ; Fix2018 . Molecular magnets with nonlocal correlations also open doors to applications in quantum information technology such as entanglement Horodecki2009 ; Bazhanov2018 ; Vargova2023 ; Bode2023 , teleportation Bennett1993 ; Krauter2013 , and quantum state transferring Lockyer2022 ; Ferrando2016 ; Soria2016 . This growing field provides diverse opportunities that harness the distinctive characteristics of molecular magnets, bridging various scientific disciplines.
@@
Quantum sensing, a key application of quantum technologies Degen2017 ; Braun2018 ; Kugelgen2021 , has witnessed advancements in various physical setups, ranging from smart materials Chugh2023 and photonic devices Pirandola2018 ; Taylor2013 to ion traps Baumgart2016 ; Bohnet2016 ; Reiter2017 ; Gilmore2021 and superconducting qubits for the purpose of quantum error correction Reed2012 ; Corcoles2015 ; Danilin2021 ; Rojkov2022 . The exploration of quantum-enhanced sensitivity is crucial in both critical and Floquet many-body quantum sensors Mishra2021 ; Mishra2022 ; Sarkar2022 ; Huang2024 . In these systems, the interaction between particles plays a central role, imparting robustness against imperfections. The study of quantum sensors aims to investigate whether projective measurements and their subsequent wave function collapse can be utilized to achieve quantum-enhanced sensitivity, opening new avenues in the field of quantum metrology Burgarth2015 ; Smerzi2018 ; Montenegro2022 ; Montenegro2023 ; Montenegro2024 ; Mihailescu2023 .
@@
Molecular magnets serve as the building blocks for a new era of customizable designer quantum sensors Troiani2019 ; Yu2021 , offering tunable functionality and adaptability to various environments Ghirri2014 . These unique capabilities pave the way for unprecedented levels of sensitivity and spatial resolution, making them useful for a broad spectrum of sensing tasks. From the study of quantum dynamic processes to the precise detection of emergent phenomena, molecular-based quantum materials that exhibit rich magnetic properties present an opportunity to control their spin states using external fields, signaling new era for future quantum devices Takahata2005 ; MolecularMagnets ; Luis2019 ; Candini2010 . Additionally, the simultaneous presence of toroidal moments in geometrically frustrated transition metal complexes with $\mathrm{C}_{3}$ symmetry presents even wider prospects, especially in the presence of strong antiferromagnetic exchange interactions Crabtree2018 ; Rao2020 ; Pister2022 . However, the practical realization of these systems as molecular quantum tools demands a deep comprehension of their electronic structure, magnetism, and their dependencies to factors such as antisymmetric contributions to magnetic exchange (Dzyalozhinsky-Moriya interaction) Robert2019 ; Liu2021 or single-ion anisotropy Rinehart2011 ; Georgiev2022 .
@@
In a recent study, a trinuclear iron(III) molecular complex $[\mathrm{Fe}_{3}\mathrm{Cl}_{3}(\mathrm{saltag}^{\mathrm{Br}})(\mathrm{py})_{6}]\mathrm{ClO}_{4}\\ \{\mathrm{H}_{5}\mathrm{saltag}^{\mathrm{Br}}=1,2,3-\mathrm{tris}[(5-\mathrm{bromo}-\mathrm{salicylidene})\mathrm{amino}]\mathrm{guanidine}\}$ (abbreviated as Fe3) with an antiferromagnetically coupled triangular structure of $\mathrm{C}_{3}$ symmetry was synthesized and characterized in detail Plass2023 . Both experimental and theoretical investigations were carried out to explore magnetic properties of this complex Plass2023 . Motivated by rich magnetic properties of the molecular compound Fe3 and rich quantum properties expected from it, in this work we consider this complex and investigate the genuine bipartite and tripartite entanglement through the behavior of negativity, a concept introduced by Peres and Horodecki Horodecki2009 ; Peres1996 ; Horodecki1996 . This analysis aims to provide insights into the intricate quantum correlations present in the molecular complex Fe3, offering a deeper understanding of its quantum properties.
Spin squeezing Kitagawa1993 ; Wineland1994 ; Wang2003 ; Wiseman2010 is a phenomenon involving the redistribution of quantum fluctuations in a minimum-uncertainty state. This phenomenon has attracted significant interest in recent years, being utilized as a tool to produce Trail2010 ; Bao2020 ; Huang2023 and detect Cheraghi2022 ; Mahdavifar2023 ; Mahdavifar2024 quantum entangled states. Most studies on spin squeezing and quantum sensing have focused on quantum spin systems with the lowest possible spin value $S=1/2$. However, exploring magnetic molecules with higher spin magnitudes could present a promising avenue for improving these properties. It has been found that increasing the number of states interacting with the magnetic field can enhance magnetic properties. Therefore, working with trinuclear high-spin iron(III) complexes composed from spin species $S=5/2$ could potentially lead to a further enhancement in quantum sensing and spin squeezing. This, in turn, would showcase the capabilities of small high-spin magnetic clusters in sensing ability and spin squeezing.
In this work, we examine the phenomenon of spin squeezing in the context of the trinuclear high-spin iron(III) complex. This study further explores spin squeezing as a tool for improving atomic precision measurements beyond the standard quantum limit within the field of quantum metrology Smerzi2018 . We discuss the theoretical understanding of spin squeezing in the molecular compound Fe3 from its creation in a molecular magnet to its proportionality with magnetic-field-driven phase transitions, where the system may act as a programmable quantum sensor highlighting its significance in advancing quantum technologies.
@@
The structure of the paper is as follows. In Sec. II we introduce the model and explore the fundamental concepts that underpin quantum sensors based on the molecular magnet Fe3. We investigate thermal bipartite and tripartite negativities of the system with respect to temperature and magnetic field in Sec. III. In Sec. IV, we discuss the results for spin squeezing and describe the behavior of the spin squeezing parameter in relation to magnetic field and temperature in the trinuclear high-spin iron(III) complex. Additionally, we investigate the quantum sensing ability of the molecular complex Fe3 using sequential measurements in Sec. V. Section VI contains a brief summary and future outlooks.
@@ II Spin-5/2 Heisenberg triangular cluster
Let us examine quantum and magnetic properties of the spin-5/2 Heisenberg triangular cluster, which represents a suitable model system for the trinuclear high-spin iron(III) molecular complex $[\mathrm{Fe}_{3}\mathrm{Cl}_{3}(\mathrm{saltag}^{\mathrm{Br}})(\mathrm{py})_{6}]\mathrm{ClO}_{4}$ referred to as Fe3. It has been convincingly evidenced in Ref. Plass2023 that the molecular compound Fe3 indeed affords, due to a high-spin state of iron(III) magnetic ions and $\mathrm{C}_{3}$ symmetry of the underlying superexchange pathways (see Fig. 1), an experimental realization of the spin-5/2 Heisenberg equilateral triangle given by the Hamiltonian:
Figure 1: Left panel displays a scheme of three exchange interactions within the trinuclear high-spin iron(III) molecular complex $[\mathrm{Fe}_{3}\mathrm{Cl}_{3}(\mathrm{saltag}^{\mathrm{Br}})(\mathrm{py})_{6}]\mathrm{ClO}_{4}$. Right panel shows the molecular structure of the same complex visualized according to crystallographic data reported in Ref. Plass2023 .
$\hat{H}$ $=$ $J\left(\hat{\bm{S}}_{1}\cdot\hat{\bm{S}}_{2}+\hat{\bm{S}}_{2}\cdot\hat{\bm{S}}_{3}+\hat{\bm{S}}_{1}\cdot\hat{\bm{S}}_{3}\right)-$ (1)
@@
Here, $\hat{S}^{\alpha}_{j}$ ($\alpha=x,y,z$) are conventional spatial components of spin-$\frac{5}{2}$ operators assigned to three iron(III) metal centers $j=1-3$ of the molecular complex Fe3, the coupling constant $J$ determines a size of intramolecular exchange interactions between the pairs of spin-5/2 iron(III) magnetic ions, $g=2.0$ is the gyromagnetic ratio of the fully isotropic spin-5/2 iron(III) magnetic ions with zero orbital angular momentum, $\mu_{B}$ is Bohr magneton and $B_{z}$ is the magnetic field applied along the $z$-direction. According to experimental results and DFT calculations using broken-symmetry states (BS-DFT) Plass2023 , this coupling constant is antiferromagnetic with a moderate value of $J=12.56\,\mathrm{cm}^{-1}$.
By introducing the total spin operator $\hat{S}_{\mathrm{T}}=\hat{S}_{1}+\hat{S}_{2}+\hat{S}_{3}$ and its $z$-component $\hat{S}_{\mathrm{T}}^{z}=\hat{S}_{1}^{z}+\hat{S}_{2}^{z}+\hat{S}_{3}^{z}$, the Hamiltonian (1) of the spin-5/2 Heisenberg triangle can be alternatively expressed within the Kambe projection method Kambe1950 ; Sinn1970 in the following form:
@@
This alternative formula provides a full set of energy eigenvalues of the spin-5/2 Heisenberg triangle and facilitates the analysis of its magnetic properties within the Kambe coupling scheme Kambe1950 ; Sinn1970 . The rising magnetic field generally causes crossing of energy levels from different sectors of $z$-component of the total spin ${S}_{\mathrm{T}}^{z}$, whereby a level-crossing field for a magnetic-field-driven transition between the sectors with ${S}_{\mathrm{T}}^{z}$ and ${S}_{\mathrm{T}}^{z}+1$ is given by the condition $g\mu_{\mathrm{B}}B^{z}_{{S}_{\mathrm{T}}^{z},{S}_{\mathrm{T}}^{z}+1}\ =(S_{\mathrm{T}}+1)J$. The precise structure of these ground states and their degeneracies are given in A.
Figure 2: The isothermal magnetization curves of the spin-5/2 Heisenberg triangle calculated for a few selected temperatures by assuming the coupling constant $J=12.56\,\mathrm{cm}^{-1}$ and the gyromagnetic ratio $g=2.0$. High-field magnetization data as reported in Ref. Plass2023 for the molecular complex Fe3 at sufficiently low temperature $T=1.8\,\mathrm{K}$ are depicted by green circles. The ground-state eigenvectors $|S_{\mathrm{T}}\rangle$ corresponding to intermediate magnetization plateaus are indicated by arrows.
@@
Figure 2 displays isothermal magnetization curves of the spin-5/2 Heisenberg triangle for a few selected temperatures including $T=1.8\,\mathrm{K}$, which corresponds to temperature of the pulsed-field magnetization measurement performed on the polycrystalline sample of the molecular complex Fe3 reported in Ref. Plass2023 . A zero-temperature magnetization curve shown in Fig. 2 by a dotted line implies presence of a sequence of intermediate magnetization plateaus at 1/15, 1/5, 1/3, 7/15, 3/5, 11/15, and 13/15 of the saturation magnetization before the magnetization reaches its saturation value. Each intermediate magnetization plateau reflects a quantum state with a given value of the $z$-component of the total spin ${S}_{\mathrm{T}}^{z}=1/2,3/2,5/2,\cdots,15/2$ and the corresponding eigenenergy (3), which becomes energetically most favorable within a specific interval of the magnetic field. The magnetization data, which were experimentally recorded for the molecular compound Fe3 at sufficiently low temperature $1.8$ K up to 60 T shown by green circles, are in a perfect agreement with the respective theoretical prediction depicted by a blue solid line. A comparison between the theoretical and experimental results thus verifies actual existence of first four quantum ground states reflected as the respective fractional magnetization plateaus at 1/15, 1/5, 1/3, and 7/15 of the saturation magnetization. As temperature increases, the intermediate magnetization plateaus gradually smear out and the magnetization curve becomes smoother.

### v3

# Multipartite Entanglement and Quantum Sensing in a Spin-5/2 Heisenberg Molecular Iron(III) Triangle

@@ Abstract
This study provides insights into the static and dynamic quantum properties of the trinuclear high-spin iron(III) molecular complex $[\mathrm{Fe}_{3}\mathrm{Cl}_{3}(\mathrm{saltag^{\mathrm{Br}}})(\mathrm{py})_{6}]\mathrm{ClO}_{4}$ to be further abbreviated as Fe3. Using exact diagonalization of a spin-5/2 Heisenberg triangle in a magnetic field, we model the corresponding quantum behavior of the molecular compound Fe3. Our rigorous analysis employs various key metrics to explore a rich quantum behavior of this molecular compound. At sufficiently low temperatures, the bipartite negativity reveals that the pairwise entanglement between any pair of iron(III) magnetic ions of the molecular complex Fe3 can be significantly enhanced by a small magnetic field. This enhancement is followed by unconventional step-like changes characterized by a sequence of plateaus and sudden downturns as the magnetic field further increases. A qualitatively similar behavior is also observed in the genuine tripartite entanglement among all three iron(III) magnetic ions in the trinuclear complex Fe3. Notably, the bipartite and tripartite entanglement persist in the molecular complex Fe3 up to moderate temperatures of approximately 30 K and 70 K, respectively. Additionally, we demonstrate the achievement of quantum-enhanced sensitivity by initializing the molecular complex Fe3 in Dicke states. Finally, we investigated a quantum-sensing protocol by applying a local magnetic field specifically to one iron(III) magnetic ion of the molecular compound Fe3 and performing readout sequentially on one of two remaining iron(III) magnetic ions.
@@ I Introduction
The field of molecular magnetism [1, 2, 3, 4] is rapidly advancing with combining insights from quantum information science, condensed matter physics and chemical engineering [6, 7, 5, 8]. Molecular magnets, with their unique features, have promising applications in ultra-high-density storage devices [9, 10], spintronics [11, 12], quantum computing [12, 13, 14, 15, 16, 17, 18], and quantum algorithms [19]. Phenomena like quantum tunneling of magnetization and molecular cooling in the quantum magnets are essential for quantum computing and achieving low temperatures [20, 21, 22, 23]. Molecular magnets with nonlocal correlations also open doors to applications in quantum information technology such as entanglement [24, 25, 26, 27], teleportation [28, 29], and quantum state transferring [30, 31, 32]. This growing field provides diverse opportunities that harness the distinctive characteristics of molecular magnets, bridging various scientific disciplines.
@@
Quantum sensing, a key application of quantum technologies [33, 34, 35], has witnessed advancements in various physical setups, ranging from smart materials [36, 37] and photonic devices [38, 39] to ion traps [40, 41, 42, 43] and superconducting qubits for the purpose of quantum error correction [44, 45, 46, 47]. The exploration of quantum-enhanced sensitivity is crucial in both critical and Floquet many-body quantum sensors [48, 50, 49, 51]. In these systems, the interaction between particles plays a central role, imparting robustness against imperfections. The study of quantum sensors aims to investigate whether projective measurements and their subsequent wave function collapse can be utilized to achieve quantum-enhanced sensitivity, opening new avenues in the field of quantum metrology [52, 53, 54, 55, 56, 57].
@@
Molecular magnets serve as the building blocks for a new era of customizable designer quantum sensors [58, 59], offering tunable functionality and adaptability to various environments [60]. These unique capabilities pave the way for unprecedented levels of sensitivity and spatial resolution, making them useful for a broad spectrum of sensing tasks. From the study of quantum dynamic processes to the precise detection of emergent phenomena, molecular-based quantum materials that exhibit rich magnetic properties present an opportunity to control their spin states using external fields, signaling new era for future quantum devices [62, 61, 63, 64]. Additionally, the simultaneous presence of toroidal moments in geometrically frustrated transition metal complexes with $\mathrm{C}_{3}$ symmetry presents even wider prospects, especially in the presence of strong antiferromagnetic exchange interactions [65, 66, 67]. However, the practical realization of these systems as molecular quantum tools demands a deep comprehension of their electronic structure, magnetism, and their dependencies to factors such as antisymmetric contributions to magnetic exchange (Dzyalozhinsky-Moriya interaction) [68, 69] or single-ion anisotropy [70, 71].
@@
In a recent study, a trinuclear iron(III) molecular complex $[\mathrm{Fe}_{3}\mathrm{Cl}_{3}(\mathrm{saltag}^{\mathrm{Br}})(\mathrm{py})_{6}]\mathrm{ClO}_{4}\\ \{\mathrm{H}_{5}\mathrm{saltag}^{\mathrm{Br}}=1,2,3-\mathrm{tris}[(5-\mathrm{bromo}-\mathrm{salicylidene})\mathrm{amino}]\mathrm{guanidine}\}$ (abbreviated as Fe3) with an antiferromagnetically coupled triangular structure of $\mathrm{C}_{3}$ symmetry was synthesized and characterized in detail [72]. Both experimental and theoretical investigations were carried out to explore magnetic properties of this complex [72]. Motivated by rich magnetic properties of the molecular compound Fe3 and rich quantum properties expected from it, in this work we consider this complex and investigate the genuine bipartite and tripartite entanglement through the behavior of negativity, a concept introduced by Peres and Horodecki [24, 73, 74]. This analysis aims to provide insights into the intricate quantum correlations present in the molecular complex Fe3, offering a deeper understanding of its quantum properties.
In this work, we examine the phenomenon of quantum entanglement in the context of the trinuclear high-spin iron(III) complex. We particularly discuss the theoretical understanding of tripartite entanglement in the molecular compound Fe3 focusing on its stepwise changes along magnetic-field-driven phase transitions. Besides, we also demonstrate quantum-enhanced sensitivity when the molecular complex Fe3 is initialized in Dicke state. A quantum-sensing protocol involves applying a local magnetic field specifically to one of the iron(III) magnetic ions in the Fe3 molecular compound and performing sequential readout on one of the two remaining iron(III) magnetic ions.
@@
The structure of the paper is as follows. In Sec. II we introduce the model and explore the fundamental concepts that underpin quantum sensors based on the molecular magnet Fe3. Then, we investigate thermal bipartite and tripartite negativities of the system with respect to temperature and magnetic field in Sec. III. In Sec. IV, the spin-5/2 Heisenberg triangle is firstly initialized in each of the four simplest Dicke states to assess its performance. Subsequently, the quantum spin dynamics are analyzed under the influence of a transverse magnetic field applied to a single subsystem of the model. Next, we investigate the quantum sensing abilities of the molecular complex Fe3 through sequential measurements. Finally, sensitivity is evaluated using the classical Fisher information (CFI) as the quantifier. Section V contains a brief summary and future outlooks.
@@ II Spin-5/2 Heisenberg triangular cluster
Let us examine quantum and magnetic properties of the spin-5/2 Heisenberg triangular cluster, which represents a suitable model system for the trinuclear high-spin iron(III) molecular complex $[\mathrm{Fe}_{3}\mathrm{Cl}_{3}(\mathrm{saltag}^{\mathrm{Br}})(\mathrm{py})_{6}]\mathrm{ClO}_{4}$ referred to as Fe3. It has been convincingly evidenced in Ref. [72] that the molecular compound Fe3 indeed affords, due to a high-spin state of iron(III) magnetic ions and $\mathrm{C}_{3}$ symmetry of the underlying superexchange pathways (see Fig. 1), an experimental realization of the spin-5/2 Heisenberg equilateral triangle given by the Hamiltonian:
Figure 1: Left panel displays a scheme of three exchange interactions within the trinuclear high-spin iron(III) molecular complex $[\mathrm{Fe}_{3}\mathrm{Cl}_{3}(\mathrm{saltag}^{\mathrm{Br}})(\mathrm{py})_{6}]\mathrm{ClO}_{4}$. Right panel shows the molecular structure of the same complex visualized according to crystallographic data reported in Ref. [72].
$\hat{H}$ $=$ $J\left(\hat{\bm{S}}_{1}\cdot\hat{\bm{S}}_{2}+\hat{\bm{S}}_{2}\cdot\hat{\bm{S}}_{3}+\hat{\bm{S}}_{1}\cdot\hat{\bm{S}}_{3}\right)$ (1)
@@
Here, $\hat{S}^{\alpha}_{j}$ ($\alpha=x,y,z$) are conventional spatial components of spin-$\frac{5}{2}$ operators assigned to three iron(III) metal centers $j=1-3$ of the molecular complex Fe3, the coupling constant $J$ determines a size of intramolecular exchange interactions between the pairs of spin-5/2 iron(III) magnetic ions, $g=2.0$ is the gyromagnetic ratio of the fully isotropic spin-5/2 iron(III) magnetic ions with zero orbital angular momentum, $\mu_{B}$ is Bohr magneton and $B_{z}$ is the magnetic field applied along the $z$-direction. According to experimental results and DFT calculations using broken-symmetry states (BS-DFT) [72], this coupling constant is antiferromagnetic with a moderate value of $J=12.56\,\mathrm{cm}^{-1}$.
By introducing the total spin operator $\hat{S}_{\mathrm{T}}=\hat{S}_{1}+\hat{S}_{2}+\hat{S}_{3}$ and its $z$-component $\hat{S}_{\mathrm{T}}^{z}=\hat{S}_{1}^{z}+\hat{S}_{2}^{z}+\hat{S}_{3}^{z}$, the Hamiltonian (1) of the spin-5/2 Heisenberg triangle can be alternatively expressed within the Kambe projection method [75, 76] in the following form:
@@
This alternative formula provides a full set of energy eigenvalues of the spin-5/2 Heisenberg triangle and facilitates the analysis of its magnetic properties within the Kambe coupling scheme [75, 76]. The rising magnetic field generally causes crossing of energy levels from different sectors of $z$-component of the total spin ${S}_{\mathrm{T}}^{z}$, whereby a level-crossing field for a magnetic-field-driven transition between the sectors with ${S}_{\mathrm{T}}^{z}$ and ${S}_{\mathrm{T}}^{z}+1$ is given by the condition $g\mu_{\mathrm{B}}B^{z}_{{S}_{\mathrm{T}}^{z},{S}_{\mathrm{T}}^{z}+1}\ =(S_{\mathrm{T}}+1)J$. The precise structure of these ground states and their degeneracies are given in A.
Figure 2: The isothermal magnetization curves of the spin-5/2 Heisenberg triangle calculated for a few selected temperatures by assuming the coupling constant $J=12.56\,\mathrm{cm}^{-1}$ and the gyromagnetic ratio $g=2.0$. High-field magnetization data as reported in Ref. [72] for the molecular complex Fe3 at sufficiently low temperature $T=1.8\,\mathrm{K}$ are depicted by green circles. The ground-state eigenvectors $|S_{\mathrm{T}}\rangle$ corresponding to intermediate magnetization plateaus are indicated by arrows.
@@
Figure 2 displays isothermal magnetization curves of the spin-5/2 Heisenberg triangle for a few selected temperatures including $T=1.8\,\mathrm{K}$, which corresponds to temperature of the pulsed-field magnetization measurement performed on the polycrystalline sample of the molecular complex Fe3 reported in Ref. [72]. A zero-temperature magnetization curve shown in Fig. 2 by a dotted line implies presence of a sequence of intermediate magnetization plateaus at 1/15, 1/5, 1/3, 7/15, 3/5, 11/15, and 13/15 of the saturation magnetization before the magnetization reaches its saturation value. Each intermediate magnetization plateau reflects a quantum state with a given value of the $z$-component of the total spin ${S}_{\mathrm{T}}^{z}=1/2,3/2,5/2,\cdots,15/2$ and the corresponding eigenenergy (3), which becomes energetically most favorable within a specific interval of the magnetic field. The magnetization data, which were experimentally recorded for the molecular compound Fe3 at sufficiently low temperature $1.8$ K up to 60 T shown by green circles, are in a perfect agreement with the respective theoretical prediction depicted by a blue solid line. A comparison between the theoretical and experimental results thus verifies actual existence of first four quantum ground states reflected as the respective fractional magnetization plateaus at 1/15, 1/5, 1/3, and 7/15 of the saturation magnetization. As temperature increases, the intermediate magnetization plateaus gradually smear out and the magnetization curve becomes smoother.